1. A certain substance was placed inside a room where the temperature is 17°C. it is observed that after 30 seconds, the temperature of the substance drops to 27°C and after 1 minute, the temperature drops to 20°C. what is the initial temperature of the body? Ans. 50.33°C
  2. The bureau of census record in 1975 shows that the population in the country doubles compared to that of 1955. In what year will the population quadruple? Ans. 1995
1
Expert's answer
2021-11-15T16:07:33-0500

"\frac{}{}" "\text{The change in temperature is given by }\\
\frac{d\theta}{dt}= k(17-\theta)\\
\implies \frac{d\theta}{17-\theta}=kdt\\
\implies \theta = 17-ce^{-kt}-(1)\\
\text{at t = 60 seconds and $\theta = 20$, we have that}\\
-3=ce^{-60k}-(3)\\
\text{at t = 30seconds and $\theta = 27$, we have that}\\
\text{Dividing (3) by (2), we have}\\
0.3 = e^{-30k}\\
\implies k = 0.0401\\
\text{Putting k =0.0401 in 3 we have that}\\
c = -33.33\\
\text{Next, we substitute k and c in (1), at time 0 to get our solution}\\
\therefore \theta = 17+33.33e^0=50.33^0C\\
\frac{dp}{dt}=kt\\
\implies \frac{dp}{p}=kt\\
\implies \ln p =kt +A\\
\implies p = Ce^{kt}\\
\text{At t = 20, p=2C, as given in the question}\\
\implies 2 = e^{20k}\\
\implies \ln 2 = 20k\\
\implies k = 0.0347\\
\text{When p=4C, we have that}\\
4 = e^{0.0347t}\\
\implies t = 39.95 \text{ approximately 40 years. Hence p quadruples at}\\
\text{1955 + 40years=1995}"